OpenStudy (anonymous):

4x^6 (3x^2 +5y^3) simplify??

OpenStudy (anonymous):

@ivettef365 ??

OpenStudy (ivettef365):

multiply the digits (4) . (3) = 12 and the exponents as follows x^6 . x^2 = x^(6+2) = x^8, so you have 12x^8 + 20x^6y^3

OpenStudy (anonymous):

thats the answer @ivettef365 ?

OpenStudy (ivettef365):

yes